OpenStudy (desiraem):

two coordinates for y= -2? help please c':

OpenStudy (chris911):

can u explain more...

OpenStudy (chris911):

@whpalmer4 can u help her

OpenStudy (desiraem):

Select the coordinates of two points on the line y = -2. a) (2, -2) , (-2, 2) b) (-2,-2) , (-2, 0) c) (2, -2) , (0, -2) d) (-2, 2) , (-2, 2) That's the question and choices.

OpenStudy (whpalmer4):

well, any value of x will do, so long as it is coupled with a y value of -2

OpenStudy (whpalmer4):

just remember that it's (x,y)

OpenStudy (desiraem):

Ugh, lol what do you mean i'm kinda slow at match x)

OpenStudy (johnweldon1993):

So with a coordinate in the form (x , y) you need 2 of them that have y = -2 so your answer will have (x , -2) (x , -2)

OpenStudy (johnweldon1993):

Note, it doesn't matter what the 'x' is...as long as it is (x , -2) and (x , -2)

OpenStudy (desiraem):

but, none of my choices have -2 as the y value?

OpenStudy (desiraem):

a) (2, -2) , (-2, 2) b) (-2,-2) , (-2, 0) c) (2, -2) , (0, -2) d) (-2, 2) , (-2, 2)

OpenStudy (johnweldon1993):

Are you sure? (x , y) ^ the -2 would be here

OpenStudy (desiraem):

omgggg, so it's c?

OpenStudy (johnweldon1993):

Yes it is C :)

OpenStudy (desiraem):

lol, sorry i'm kinda slow (x and thanks (:

OpenStudy (johnweldon1993):

Don't worry about it...you still got it :)

OpenStudy (chris911):

thanks @johnweldon1993 and @whpalmer4 for helping her out :)

OpenStudy (desiraem):

thanks chriss! (:

OpenStudy (chris911):

no prob ;)

OpenStudy (desiraem):

wait, can you guys help me with another one?

OpenStudy (desiraem):

Write the equation of the line that is perpendicular to the line y = \[\frac{ 3 }{ 2}\]x + 4 and passes through the point (-6, 3 ) A) Y=\[\frac{ 3 }{ 2 }x - 1\]

OpenStudy (desiraem):

B) \[\frac{ 3 }{ 2 }x +3\] C) Y= \[-\frac{ 2 }{ 3 }x -1\] D) Y= \[-\frac{ 2 }{ 3 }x + 3\]